OpenStudy (anonymous):

Solve the Equation!!!!!!! PLEASE HELP! 3x^2=x-9

OpenStudy (kc_kennylau):

There are three methods to deal with quadratic equations (equation in the second degree) like this: (a) Cross method (b) Complete the square (c) Quadratic formula

OpenStudy (kc_kennylau):

Please choose one.

OpenStudy (tkhunny):

...among others.

OpenStudy (anonymous):

i guess whichever is easier

OpenStudy (tkhunny):

Bad idea. In school - Whichever will teach you the most. In general - Whichever works the best and can be explained. In programming - Which is fastest or most appropriate or more effective at eliminating round-off error.

OpenStudy (anonymous):

uhh can i just get some help?

OpenStudy (anonymous):

i need to go to bed sometime lol

OpenStudy (anonymous):

http://www.slideshare.net/leblance/chapter-4-extra-practice-answers Answer sheet..number 79

OpenStudy (tkhunny):

You've received quite a lot. What method do you choose? Show us!

OpenStudy (anonymous):

ive always thought u factored

OpenStudy (tkhunny):

Can you factor it? Show that, please.

OpenStudy (anonymous):

i dont know if i can, all i have so far is probably wrong. I need someone that knows what their doing

OpenStudy (anonymous):



OpenStudy (kc_kennylau):

Don't divide all by -3 because that'll make the middle coefficient non-integer

OpenStudy (anonymous):

well then can u show me the full problem step by step so i can understand???

OpenStudy (kc_kennylau):

Wait, this equation doesn't have real roots

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

look at answer key

OpenStudy (kc_kennylau):

Are you learning complex number?

OpenStudy (kc_kennylau):

Oh that's why

OpenStudy (anonymous):

yup

OpenStudy (kc_kennylau):

So quadratic formula is the way

OpenStudy (anonymous):

how! show work ive already tried everything under the sun!

OpenStudy (kc_kennylau):

Quadratic formula: For ANY quadratic equation \(\Large ax^2+bx+c=0\): \[\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (kc_kennylau):

3x^2=x-9 3x^2-x+9=0 x=(1±sqrt(1-4*3*9))/6

OpenStudy (kc_kennylau):

Simplify everything

OpenStudy (anonymous):

----> http://prntscr.com/3b20m4

OpenStudy (anonymous):

@SOS101 i know thats the answer but, how do u do it? Does that website show work?

OpenStudy (anonymous):

No it doesn't ,Let me try the other website

OpenStudy (anonymous):

ok thanks @SOS101

OpenStudy (kc_kennylau):

Would you guys please stop depending on computer and learn the knowledge

OpenStudy (kc_kennylau):

Which part of my solution did you not understand

OpenStudy (anonymous):

i got down to x=1 +/- sq rt -108 over 6.......and without the internet i wouldnt know how to do half the stuf iknow how to do. so..

OpenStudy (anonymous):

Dude stop being so butt hurt , we live in a mobile/technology generation where if we some how need a little extra help then we can turn to technology , there's nothing wrong with receiving help

OpenStudy (kc_kennylau):

There's nothing wrong with learning things from the Internet, but absolutely with finding answers from the internet

OpenStudy (kc_kennylau):

without any explanation at all

OpenStudy (kc_kennylau):

What can you learn from your website

OpenStudy (anonymous):

Nope that's still only your opinion

OpenStudy (anonymous):

Actually I was trying to find the answer with the explanation so..

OpenStudy (anonymous):

if u read above, i said i wanted an explanation with the answer to TEACH me so..

OpenStudy (kc_kennylau):

@ShelbyRenaebb: Anything you still don't understand?

OpenStudy (anonymous):

btw i got the answer, using a different way. can you help with Number 37 & 41??? i really need to get to bed im so tired its 4 am D:

OpenStudy (kc_kennylau):

It's still the quadratic formula

OpenStudy (kc_kennylau):

For 41 it's not the quadratic formula

OpenStudy (anonymous):

can u teach me how to graph them?? i dont know how and want to learn

OpenStudy (kc_kennylau):

\[\begin{array}{rcl} x^2+4x&=&0\\ x(x+4)&=&0\\ x=0&or&x+4=0\\ x=0&or&x=-4 \end{array}\]

OpenStudy (anonymous):

my test is in 2 days

OpenStudy (kc_kennylau):

Basically you can just find y when x=-2, x=-1, x=0, x=1, and x=2 then join the dots together if you wanna graph 'em

OpenStudy (anonymous):

how do u find the answer.. (#37)? its supposed to be 0.41 and -2.41

OpenStudy (kc_kennylau):

For 37 it's just the quadratic equation

OpenStudy (anonymous):

kk thanks